OpenStudy (anonymous):

A rectangle has an area of 16 square feet. Its length and width are whole numbers. What is its maximum perimeter? 32 feet 64 feet 34 feet 16 feet

OpenStudy (campbell_st):

so list the factors of 16 the perimeter is P = 2* factor 1 + 2* factor 2 this will allow you to find the max perimeter

OpenStudy (anonymous):

you lost me, im a visual learner will you explain it some more @campbell_st

OpenStudy (campbell_st):

ok... what whole numbers multiply to 16 can you list them in pairs..?

OpenStudy (anonymous):

1*16 8*2

OpenStudy (campbell_st):

and there is also 4 * 4 so perimeter = 2* length + 2 * width so using your measurements above 1* 16 P = 2*1 + 2*16 2*8 P = 2*2 + 2*8 4*4 P = 2*4 + 2*4 just evaluate to find the rectangle with the largest perimeter.

OpenStudy (anonymous):

it cant be4*4 that would be a square @campbell_st

OpenStudy (campbell_st):

well a square is a specific rectangle.. but it has the smallest perimeter anyway

OpenStudy (anonymous):

34 is the biggest right?

OpenStudy (campbell_st):

yes thats it